Shuki:

Algebra 2 math. Can someone help me with this, I'm having a hard time understanding and solving these equations.

Laylalyssa:

yea i might be able to help, post ur question

Shuki:

Laylalyssa:

for substitution u jus gotta solve on equation for one variable, then substitute the variable into the other question for the other variable

563blackghost:

For the first one, you should make the equation to that of y or x so you can input it into the 2nd equation. Can you simplify the question to that of y?

Laylalyssa:

u still there ?

Shuki:

yes im just trying to write the equations down

Shuki:

im still really confused im sorry x-y=2 and 2x+3y=9, do i put them together?

Shuki:

if so i have 3x+4y=11

Laylalyssa:

nah, good try but no u have to pick a variable to solve for, and then choose an equation to solve the variable for

Shuki:

Oh right, it doesnt matter which equation i choose first right?

Laylalyssa:

nah, u can choose whichever one looks easier to you

Laylalyssa:

Once u got the variable for one of the equations, then u plug it into the other equation, and solve. So for example, if y was equal to 7, then you would go into the other equation and plug in "7" everywhere that you see "y", and then solve for x

Laylalyssa:

I'm not saying 'y' is equal to 7 tho, thats jus an example u get it now?

Shuki:

i understood the plug in part, just a bit stuck on finding the variable

Laylalyssa:

oh aright, so basically, you know how to solve an equation right

Laylalyssa:

If ur solving for X, you gotta isolate x, and get it on one side by itself

Shuki:

yeah i got that now

Shuki:

thank you

Laylalyssa:

yw